# Quantum algorithm for partial differential equations of non-conservative systems with spatially varying parameters

### v1

@@ I Introduction
Quantum computing emerges as a promising alternative that offers the potential to perform more efficient computation than classical computing, leveraging the principles of quantum mechanics. There are several potential applications, including quantum phase estimation [7], linear system solvers [8, 9, 10], and Hamiltonian simulation [11, 12, 7, 13, 14, 15]. Hamiltonian simulation is a technique to simulate a time evolution of a quantum system with the Hamiltonian $\mathcal{H}$; its aim is to implement the time evolution operator $\exp(-i\mathcal{H}\tau)$ with the time increment $\tau$ on a quantum computer. The typical system Hamiltonian includes Ising Hamiltonians [13, 14] and molecular Hamiltonians [15]. Several works also applied Hamiltonian simulation techniques for simulating the classical conservative systems by mapping the governing equations of classical systems to the Schrödinger equation [16, 17, 18]. Costa et al. [16] proposed a method of simulating the wave equation via Hamiltonian simulation where the system Hamiltonian is given as the incidence matrix of the discretized simulation space. Babbush et al. [17] extended this approach to propose a quantum simulation technique to simulate classical coupled oscillators with the rigorous proof of the exponential speedup over any classical algorithms. Miyamoto et al. [18] proposed a Hamiltonian-simulation-based quantum algorithm that solves the Vlasov equation for large-scale structure simulation. Recently, Schade et al. [19] posed a quantum computing concept for simulating one-dimensional elastic wave propagation in spatially varying density and provide the potential of exponential speedup. Note that these approaches are limited to handling PDEs of conservative systems that can be mapped to the Schrödinger equation; that is, the system with time evolution operator represented by a unitary operator. Because most classical systems are non-conservative, it is important to develop a Hamiltonian-simulation-based approaches applicable to those practical systems.
@@
The rest of this paper is organized as follows. In Sec. II, we briefly introduce the LCHS method [25] and the finite difference operators along with its representation in a qubit system. In Sec. III, we first define the problem of interest, followed by its discretization onto the qubit system. In particular, we will describe how the Hamiltonian describing the spatially varying material properties can be constructed in an efficient manner using a logic minimization technique. We also give a tensor-network-based technique for approximately constructing the quantum circuit for LCU. In Sec. IV, we provide several numerical experiments, applying the proposed method to the acoustic and heat equations. Finally, we conclude this paper in Sec. V.
@@ II Preliminaries
Section II A briefly reviews the method of linear combination of Hamiltonian simulation (LCHS) [25] for representing a non-conservative dynamics in non-unitary quantum mechanical framework. Section II B summarizes the overview of the proposed method, describing what we tackle to apply LCHS for PDEs of non-conservative systems with spatially varying parameters. Section II C introduces the finite difference operators which will be used for mapping a classical PDE to a quantum non-unitary dynamics.
@@ II.1 Linear combination of Hamiltonian simulation
    Prepare $n$ qubits (referred to as the system register) to embed the solution $\bm{w}(t)$ of the ODE and $n_{\mathrm{a}}$ ancilla qubits (referred to as the ancilla register) for the LCU. The number of necessary ancilla qubits depends on discretization of $k$, that is, the number of integration points $M$ in the numerical integration with respect to $k$, in the form $n_{\mathrm{a}}=\lceil\log_{2}M\rceil$. With the integration points $\{k_{a}\}_{a=0}^{M-1}$, we let $c_{a}$ denote the coefficient of LCU, i.e., $c_{a}:=\omega_{a}/\pi(1+k_{a}^{2})$ where $\omega_{a}$ is the weight for numerical integration, and approximate Eq. (3) by $\bm{w}(t)\approx\sum_{a=0}^{M-1}c_{a}\exp\left(-i\left(\bm{H}+k_{a}\bm{L}\right)t\right)\bm{w}(0)$. Since $U_{a}:=\exp\left(-i\left(\bm{H}+k_{a}\bm{L}\right)t\right)$ is a unitary, this is exactly LCU of $\sum_{a=0}^{M-1}c_{a}U_{a}$.
@@
    Initialize the qubits to |$\ket{0}^{\otimes n_{\mathrm{a}}}\otimes\ket{0}^{\otimes n}$.
@@
Following this procedure, we obtain the output state of whole system before the measurement as $(1/\|\bm{c}\|_{1})\ket{0}^{\otimes n_{\mathrm{a}}}\otimes\exp\left(-\bm{A}t\right)\ket{w(0)}+\ket{\perp}$, where $\ket{\perp}$ represents a state orthogonal to the first term. Then, when the measurement result of the ancilla register is all zeros, the system register approximates the state (3) in the sense of numerical integration (i.e., approximation via quadrature by parts and Trotterization). The expected number of repetition $r$ to successfully get the measurement outcome of all zeros in the ancilla register is of the order $\mathcal{O}(\|\bm{c}\|_{1}^{2}\|\bm{w}(0)\|^{2}/\|\bm{w}(T)\|^{2})$, but this can be improved to $\mathcal{O}(\|\bm{c}\|_{1}\|\bm{w}(0)\|/\|\bm{w}(T)\|)$ using the amplitude amplification technique (note that $\|\bm{w}(0)\|\geq\|\bm{w}(T)\|$ in general).
II.2 Overview of the proposed method
In the procedure of LCHS, the required oracles are the state preparation oracle $O_{\mathrm{prep}}$, the coefficient oracle for the LCU $O_{\mathrm{coef}}$, the select oracle $\texttt{SEL}_{L}(s)$, and the Hamiltonian simulation oracle $O_{H}(s)$. The state preparation oracle $O_{\mathrm{prep}}$ can be implemented naively for simple initial states, and even for general cases, methods for approximate implementation using tensor networks or parameterized quantum circuits are known [26, 27, 28]. The Hamiltonian simulation oracle with finite difference operators can be efficiently implemented as described in the previous method [29]. In the same manner, the select oracle $\texttt{SEL}_{L}(s)$ can be efficiently implemented. Here, efficiently means that they can be implemented with polynomial-depth quantum circuits. However, the previous method focused on PDEs with spatially uniform parameters. Thus, the present study proposes the treatment of the spatially varying parameters of PDEs in Section III.2. Furthermore, we will consider the implementation of the coefficient oracle for the LCU, $O_{\mathrm{coef}}$ in Section III.3. Figure 1 shows an overview of our proposed method, which will be described in detail with the concrete algorithm in Section III.4.
Figure 1: Overview of the proposed method. (a) Quantum circuit of LCHS based on the work by An et al. [25]. (b) Efficient representation of Hamiltonian via logic minimizer techniques, which is described in Section III.2. (c) Implementation of the coefficient oracle for LCU, $O_{\mathrm{coef}}$, via tensor network techniques, which is described in Section III.3. (d) Quantum circuit of the select oracle $\texttt{SEL}_{L}$, which is described in Section III.4.
@@ II.2 Overview of the proposed method
II.3 Finite difference operators and its representation in qubits
Let us consider a $d$-dimensional domain $\Omega\subset\mathbb{R}^{d}$ with an orthogonal coordinate $\bm{x}:=(x_{0},\dots x_{d-1})$ and a scalar field $u(\bm{x})$ defined on the domain. Discretizing the domain $\Omega$ into a grid consisting of $N$ nodes, we represent the scalar field $u(\bm{x})$ as the value at each node, $\bm{u}:=[u(\bm{x}^{[0]}),u(\bm{x}^{[1]}),\dots,u(\bm{x}^{[N-1]})]^{\top}$, where $\bm{x}^{[j]}$ is the coordinate of the $j$-th node. The finite difference method (FDM) approximates the spatial gradient of the scalar field using its node values $\bm{u}$. For instance, the forward difference scheme approximates the first-order spatial derivative as
$\frac{\partial u(\bm{x}^{[j]})}{\partial x_{\mu}}=\frac{u(\bm{x}^{[j^{\prime}]})-u(\bm{x}^{[j]})}{h}+O(h),$ (5)
where $u(\bm{x}^{[j^{\prime}]})=u(\bm{x}^{[j]}+h\bm{e}_{\mu})$ with the orthonormal basis $\bm{e}^{[\mu]}$ along with the $x_{\mu}$-axis and $h$ is the interval of adjacent nodes. That is, $\bm{x}^{[j]}+h\bm{e}_{\mu}$ corresponds to the coordinate of the $j^{\prime}$-th node which locates next to the $j$-th node along with the $x_{\mu}$-axis. In the previous research [29], we showed that the finite difference operators including forward, backward and central differencing schemes, can be represented using operators for qubit systems, which we will briefly review here.
To represent the finite difference operator in terms of qubit operators, we first map the discretized scalar field $\bm{u}$ into the system state, with the assumption that $N=2^{n}$, as
@@ II.3 Finite difference operators and its representation in qubits
where |$\ket{j}:=\ket{j_{n-1}j_{n-2}\dots j_{0}}$ with $j_{n-1},j_{n-2},\dots,j_{0}\in\{0,1\}$ is the computational basis. Let $\sigma_{01}$ and $\sigma_{10}$ denote the ladder operators defined as
@@
Then, the shift operators $S^{-}=\sum_{j=1}^{2^{n}-1}\ket{j-1}\!\bra{j}$ | and $S^{+}=\sum_{j=1}^{2^{n}-1}\ket{j}\!\bra{j-1}$ | acting on the system are represented as
$S^{-}$ $:=\sum_{j=1}^{n}I^{\otimes(n-j)}\otimes\sigma_{01}\otimes\sigma_{10}^{\otimes(j-1)}$
$=\sum_{j=1}^{n}s^{-}_{j},$ (8)
$S^{+}$ $:=\left(S^{-}\right)^{\dagger}$
$=\sum_{j=1}^{n}I^{\otimes(n-j)}\otimes\sigma_{10}\otimes\sigma_{01}^{\otimes(j-1)}$
$=\sum_{j=1}^{n}s^{+}_{j},$ (9)
@@
$s^{-}_{j}$ $:=I^{\otimes(n-j)}\otimes\sigma_{01}\otimes\sigma_{10}^{\otimes(j-1)},$ (10)
$s^{+}_{j}$ $:=I^{\otimes(n-j)}\otimes\sigma_{10}\otimes\sigma_{01}^{\otimes(j-1)}.$ (11)
@@
$\frac{1}{h}\left(S^{-}-I^{\otimes n}\right)\sum_{j=0}^{2^{n}-1}u_{j}\ket{j}=\sum_{j=0}^{2^{n}-1}\frac{u_{j+1}-u_{j}}{h}\ket{j},$ (12)
where $u_{2^{n}}=0$, which exactly corresponds to the forward difference operator with the Dirichlet boundary condition.

### v2

@@ I Introduction
Quantum computing emerges as a promising alternative that offers the potential to perform more efficient computation than classical computing, leveraging the principles of quantum mechanics. There are several potential applications, including quantum phase estimation [7], linear system solvers [8, 9, 10], and Hamiltonian simulation [11, 12, 7, 13, 14, 15]. Hamiltonian simulation is a technique to simulate a time evolution of a quantum system with the Hamiltonian $\mathcal{H}$; its aim is to implement the time evolution operator $\exp(-i\mathcal{H}\tau)$ with the time increment $\tau$ on a quantum computer. The typical system Hamiltonian includes Ising Hamiltonians [13, 14] and molecular Hamiltonians [15]. Several works also applied Hamiltonian simulation techniques for simulating the classical conservative systems by mapping the governing equations of classical systems to the Schrödinger equation [16, 17, 18, 19]. Costa et al. [16] proposed a method of simulating the wave equation via Hamiltonian simulation where the system Hamiltonian is given as the incidence matrix of the discretized simulation space. Babbush et al. [17] extended this approach to propose a quantum simulation technique to simulate classical coupled oscillators with the rigorous proof of the exponential speedup over any classical algorithms. Miyamoto et al. [18] proposed a Hamiltonian-simulation-based quantum algorithm that solves the Vlasov equation for large-scale structure simulation. Wright et al. [19] simulated one-dimensional wave equation using a real quantum device. Recently, Schade et al. [20] posed a quantum computing concept for simulating one-dimensional elastic wave propagation in spatially varying density and provide the potential of exponential speedup. Note that these approaches are limited to handling PDEs of conservative systems that can be mapped to the Schrödinger equation; that is, the system with time evolution operator represented by a unitary operator. Because most classical systems are non-conservative, it is important to develop a Hamiltonian-simulation-based approaches applicable to those practical systems.
@@
Figure 1 shows an overview of our proposed method, which will be described in detail with the concrete algorithm in Section III.4. Our algorithm is based on LCHS, which will be reviewed in Section II.1. The quantum circuit for LCHS is summarized in Fig. 1(a), which requires the state preparation oracle $O_{\mathrm{prep}}$, the coefficient oracle for the LCU $O_{\mathrm{coef}}$, the Hamiltonian simulation oracle $O_{H}(\tau)$, and the select oracle $\texttt{SEL}_{L}(\tau)$. The state preparation oracle $O_{\mathrm{prep}}$ can be implemented naively for simple initial states, and even for general cases, methods for approximate implementation using tensor networks or parameterized quantum circuits are known [27, 28, 29]. The Hamiltonian simulation oracle with finite difference operators can be efficiently implemented as described in the previous method [30]. In the same manner, the select oracle $\texttt{SEL}_{L}(\tau)$ can be efficiently implemented. Here, efficiently means that they can be implemented with polynomial-depth quantum circuits. However, the previous method focused on PDEs with spatially uniform parameters. Thus, the present study proposes the method for simulating PDEs of non-conservative systems with spatially-varying parameters, in Section III.2. In particular, we will describe how the Hamiltonian describing the spatially varying parameters can be constructed in an efficient manner using a logic minimization technique (Fig. 1(b)). Furthermore, we will consider the approximate implementation of the coefficient oracle for the LCU, $O_{\mathrm{coef}}$, based on a tensor network technique in Section III.3 (Fig. 1(c)).
Based on the notation used in Section. II, our algorithm requires $n_{\mathrm{anc}}\sim\mathcal{O}(\log(\|\bm{L}\|_{2}T/\varepsilon^{2}))$ ancilla qubits where $\bm{L}$ is the Hermitian part of the coefficient matrix, $T$ is the simulation time, $\varepsilon$ is the precision parameter. Owing to the tensor-network-based approximation, we can implement the coefficient oracle $O_{\mathrm{coef}}$ using $\mathcal{O}(n_{\mathrm{anc}})$ non-local gates. Assuming that the Hamiltonian including spatially varying parameters is represented by at most $\mathrm{poly}(n)$ terms where the space is discretized into $2^{n}$ nodes, we can implement the Hamiltonian simulation oracle $O_{H}(\tau)$ using at most $\mathcal{O}(\mathrm{poly}(n)T^{3/2}/\varepsilon_{t}^{1/2})$ non-local gates [30] where $\varepsilon_{t}$ is the precision parameter of the time integration. Similarly, we can implement the select oracle $\texttt{SEL}_{L}(\tau)$ using at most $\mathcal{O}(n_{\mathrm{anc}}\mathrm{poly}(n)T^{3/2}/\varepsilon_{t}^{1/2})$ non-local gates. Note that we herein assume that these oracles are implemented by the second-order Suzuki-Trotter decomposition while the state-of-the-art algorithms [31, 32] achieved the time complexity $\mathcal{O}(\mathrm{poly}(\mathrm{log}(1/\varepsilon_{t})))$. The expected number of repetition to measure the outcome of all zeros in the ancilla qubits, which is success, is of the order $\mathcal{O}(\|\bm{w}(0)\|/\|\bm{w}(T)\|)$, which is the ratio of the norm of the state vector at the initial and final times.
Figure 1: Overview of the proposed method. (a) Quantum circuit of LCHS based on the work by An et al. [26]. (b) Efficient representation of Hamiltonian via logic minimizer techniques, which is described in Section III.2. (c) Implementation of the coefficient oracle for LCU, $O_{\mathrm{coef}}$, via the tensor network techniques, which is described in Section III.3. (d) Quantum circuit of the select oracle $\texttt{SEL}_{L}$, which is described in Section III.4.
@@ II Preliminaries
Section II.1 briefly reviews the method of linear combination of Hamiltonian simulation (LCHS) [26] for representing a non-conservative dynamics in non-unitary quantum mechanical framework. Section II.2 introduces the finite difference operators which will be used for mapping a classical PDE to a quantum non-unitary dynamics.
@@ II.1 Linear combination of Hamiltonian simulation
    Prepare $n$ qubits (referred to as the system register) to embed the solution $\bm{w}(t)$ of the ODE and $n_{\mathrm{anc}}$ ancilla qubits (referred to as the ancilla register) for the LCU. The number of necessary ancilla qubits depends on discretization of $k$, that is, the number of integration points $M$ in the numerical integration with respect to $k$, in the form $n_{\mathrm{anc}}=\lceil\log_{2}M\rceil$. With the integration points $\{k_{a}\}_{a=0}^{M-1}$, we let $c_{a}$ denote the coefficient of LCU, i.e., $c_{a}:=\omega_{a}/\pi(1+k_{a}^{2})$ where $\omega_{a}$ is the weight for numerical integration, and approximate Eq. (3) by $\bm{w}(t)\approx\sum_{a=0}^{M-1}c_{a}\exp\left(-i\left(\bm{H}+k_{a}\bm{L}\right)t\right)\bm{w}(0)$. Since $U_{a}:=\exp\left(-i\left(\bm{H}+k_{a}\bm{L}\right)t\right)$ is a unitary, this is exactly LCU of $\sum_{a=0}^{M-1}c_{a}U_{a}$.
@@
    Initialize the qubits to |$\ket{0}^{\otimes n_{\mathrm{anc}}}\otimes\ket{0}^{\otimes n}$.
@@
Following this procedure, we obtain the output state of whole system before the measurement as $(1/\|\bm{c}\|_{1})\ket{0}^{\otimes n_{\mathrm{anc}}}\otimes\exp\left(-\bm{A}t\right)\ket{w(0)}+\ket{\perp}$, where $\ket{\perp}$ represents a state orthogonal to the first term. Then, when the measurement result of the ancilla register is all zeros, the system register approximates the state (3) in the sense of numerical integration (i.e., approximation via quadrature by parts and Trotterization). The expected number of repetition to successfully get the measurement outcome of all zeros in the ancilla register is of the order $\mathcal{O}(\|\bm{c}\|_{1}^{2}\|\bm{w}(0)\|^{2}/\|\bm{w}(T)\|^{2})$, but this can be improved to $\mathcal{O}(\|\bm{c}\|_{1}\|\bm{w}(0)\|/\|\bm{w}(T)\|)$ using the amplitude amplification technique (note that $\|\bm{w}(0)\|\geq\|\bm{w}(T)\|$ in general).
@@
II.2 Finite difference operators and its representation in qubits
Let us consider a $d$-dimensional domain $\Omega\subset\mathbb{R}^{d}$ with an orthogonal coordinate $\bm{x}:=(x_{0},\dots x_{d-1})$ and a scalar field $u(\bm{x})$ defined on the domain. Discretizing the domain $\Omega$ into a grid consisting of $N=\prod_{\mu=0}^{d-1}N_{\mu}$ nodes, we represent the scalar field $u(\bm{x})$ as the value at each node, $\bm{u}:=[u(\bm{x}^{[0]}),u(\bm{x}^{[1]}),\dots,u(\bm{x}^{[N-1]})]^{\top}$, where $\bm{x}^{[j]}$ is the coordinate of the $j$-th node. The finite difference method (FDM) approximates the spatial gradient of the scalar field using its node values $\bm{u}$. For instance, the forward difference scheme approximates the first-order spatial derivative as
$\frac{\partial u(\bm{x}^{[j]})}{\partial x_{\mu}}=\frac{u_{j_{(\mu,1)}}-u_{j}}{h}+O(h),$ (5)
where $u_{j_{(\mu,1)}}:=u(\bm{x}^{[j]}+h\bm{e}_{\mu})$ with the orthonormal basis $\bm{e}_{\mu}$ along with the $x_{\mu}$-axis, $u_{j}:=u(\bm{x}^{[j]})$ and $h$ is the interval of adjacent nodes. That is, $\bm{x}^{[j]}+h\bm{e}_{\mu}$ corresponds to the coordinate of the $j_{(\mu,1)}$-th node which locates next to the $j$-th node along with the $x_{\mu}$-axis. Similarly, the central difference scheme with the fourth-order accuracy is given as
$\frac{\partial u(\bm{x}^{[j]})}{\partial x_{\mu}}=\frac{-u_{j_{(\mu,2)}}+8u_{j_{(\mu,1)}}-8u_{j_{(\mu,-1)}}+u_{j_{(\mu,-2)}}}{12h}+O(h^{4}),$ (6)
where $u_{j_{(\mu,2)}}:=u(\bm{x}^{[j]}+2h\bm{e}_{\mu})$, $u_{j_{(\mu,-1)}}:=u(\bm{x}^{[j]}-h\bm{e}_{\mu})$ and $u_{j_{(\mu,-2)}}:=u(\bm{x}^{[j]}-2h\bm{e}_{\mu})$. In the previous research [30], we showed that the finite difference operators including forward, backward and central differencing schemes, can be represented using operators for qubit systems, which we will briefly review here.
To represent the finite difference operator in terms of qubit operators, we first map the discretized scalar field $\bm{u}$ into the system state, with the assumption that $N_{\mu}=2^{n_{\mu}}$, as
@@ II.2 Finite difference operators and its representation in qubits
where $n:=\sum_{\mu=0}^{d-1}n_{\mu}$, |$\ket{j}:=\ket{j_{n-1}j_{n-2}\dots j_{0}}$ with $j_{n-1},j_{n-2},\dots,j_{0}\in\{0,1\}$ is the computational basis. Let $\sigma_{01}$ and $\sigma_{10}$ denote the ladder operators defined as
@@
Then, the shift operators $S^{-}(n_{\mu}):=\sum_{j^{\prime}=1}^{2^{n_{\mu}}-1}\ket{j^{\prime}-1}\!\bra{j^{\prime}}$ | and $S^{+}(n_{\mu}):=\sum_{j^{\prime}=1}^{2^{n_{\mu}}-1}\ket{j^{\prime}}\!\bra{j^{\prime}-1}$ | acting on the system are represented as
$S^{-}(n_{\mu})$ $=\sum_{j^{\prime}=1}^{n}I^{\otimes(n_{\mu}-j^{\prime})}\otimes\sigma_{01}\otimes\sigma_{10}^{\otimes(j^{\prime}-1)}$
$=\sum_{j^{\prime}=1}^{n_{\mu}}s^{-}(n_{\mu},j^{\prime}),$ (9)
$S^{+}(n_{\mu})$ $=\left(S^{-}(n_{\mu})\right)^{\dagger}$
$=\sum_{j^{\prime}=1}^{n_{\mu}}I^{\otimes(n_{\mu}-j^{\prime})}\otimes\sigma_{10}\otimes\sigma_{01}^{\otimes(j^{\prime}-1)}$
$=\sum_{j^{\prime}=1}^{n_{\mu}}s^{+}(n_{\mu},j^{\prime}),$ (10)
@@
$s^{-}(n_{\mu},j^{\prime})$ $:=I^{\otimes(n_{\mu}-j^{\prime})}\otimes\sigma_{01}\otimes\sigma_{10}^{\otimes(j^{\prime}-1)},$ (11)
$s^{+}(n_{\mu},j^{\prime})$ $:=I^{\otimes(n_{\mu}-j^{\prime})}\otimes\sigma_{10}\otimes\sigma_{01}^{\otimes(j^{\prime}-1)}.$ (12)
@@
$\frac{1}{h}\left(S^{-}(n_{\mu};\{n_{\mu}\})-I^{\otimes n}\right)\sum_{j=0}^{2^{n}-1}u(\bm{x}^{[j]})\ket{j}=\sum_{j=0}^{2^{n}-1}\frac{u_{j_{(\mu,1)}}-u_{j}}{h}\ket{j},$ (13)
where $S^{-}(n_{\mu};\{n_{\mu}\}):=I^{\otimes\sum_{\mu^{\prime}=\mu+1}^{d-1}n_{\mu^{\prime}}}\otimes S^{-}(n_{\mu})\otimes I^{\otimes\sum_{\mu^{\prime}=0}^{\mu-1}n_{\mu^{\prime}}}$, and $u_{j_{(\mu,1)}}=0$ when $\bm{x}^{[j]}+h\bm{e}_{\mu}$ is out of the domain $\Omega$, which exactly corresponds to the forward difference operator with the Dirichlet boundary condition. Similarly, we obtain the following relationship for the central difference scheme with the fourth-order accuracy:
$\frac{1}{12h}\left(-S^{-}(n_{\mu};\{n_{\mu}\})^{2}+8S^{-}(n_{\mu};\{n_{\mu}\})-8S^{+}(n_{\mu};\{n_{\mu}\})+S^{+}(n_{\mu};\{n_{\mu}\})^{2}\right)\sum_{j=0}^{2^{n}-1}u(\bm{x}^{[j]})\ket{j}$
$=\sum_{j=0}^{2^{n}-1}\frac{-u_{j_{(\mu,2)}}+8u_{j_{(\mu,1)}}-8u_{j_{(\mu,-1)}}+u_{j_{(\mu,-2)}}}{12h}\ket{j},$ (14)
where $u_{j_{(\mu,2)}}$, $u_{j_{(\mu,1)}}$, $u_{j_{(\mu,-1)}}$ and $u_{j_{(\mu,-2)}}$ are zero, respectively when the nodes on $\bm{x}^{[j]}+2h\bm{e}_{\mu}$, $\bm{x}^{[j]}+h\bm{e}_{\mu}$, $\bm{x}^{[j]}-h\bm{e}_{\mu}$, and $\bm{x}^{[j]}-2h\bm{e}_{\mu}$ are out of the domain $\Omega$, which exactly corresponds to the central difference operator with the Dirichlet boundary condition.